A dense plastic toy of mass 2.4 kg is floating just beneath the surface of a pond. What is the buoyant force on it?

You lower a 2.8 kg solid gold statue into a container of water and measure the volume of displaced water. What volume will verify that it is pure gold if the density of gold is 19.3 g/cm3?

A 4 meter-long hose of 7 cm diameter is con- nected to a faucet, and a 3 meter-long hose, which has a diameter of 14 cm, is connected to the end of the first hose. At the open end of the second hose water flows out at a rate of 4 liters/minute. What is the ratio of the speed of the water in the second hose to the speed of the water flowing in the first hose?
Answer #1

Buoyant force Fb = mg

Fb = 2.4 * 9.81

Fb = 23.544 N

-------------------------


density = mass/volume

Desnity of gold = 19.3g/cm^3 = 19300 kg/m^3

so

Volume = m/D

V = 2.8/(19300)

V = 1.45*10^-4 m^3

-----------------